OpenStudy (anonymous):

10. If F(x) is the function f(g(h(x))) and F(x) = (4+3sqrt x)^9 , list the functions for each of the following: f(x), g(x) and h(x). Specify the domain and the range for F(x).

OpenStudy (anonymous):

so we have to split F(x) into 3 functions. i guess we can do it this way: h(x) = sqrt(x) g(x) = 4+3x f(x) = x^9 so f(g(h(x))) = f(g(sqrt(x))) = f(4+3sqrt(x)) = (4+3sqrt(x))^9 = F(x) clear?

OpenStudy (anonymous):

Thank you that was very clear!

OpenStudy (anonymous):

:)